OpenStudy (awkwardpanda):

For each function, find f(-2), f(-0.5), and f(3). 1. f(x)=-x+4 2. f(x)= 3/8x-3

OpenStudy (anonymous):

f(-2)=-2+4, f(-0.5)=-0.5+4, f(3)=-3+4

OpenStudy (shamil98):

f(x) = -(-2) +4 f(x) = -(0.5) + 4 f(x) = -(3) +4

OpenStudy (shamil98):

Justin you forgot that x is multiplied with -1. :)

OpenStudy (anonymous):

f(-2)=3/8*-2-3, f(-0.5)=3/8*-0.5-3, f(3)=3/8*3-3

OpenStudy (anonymous):

dammit shamil. xD sorry

OpenStudy (shamil98):

for the second one. f(x) = (3/8)(-2) -3 f(x) = (3/8)(0.5) -3 f(x)= (3/8)(3) -3

OpenStudy (shamil98):

its not 3/8x it's (3/8)x

OpenStudy (awkwardpanda):

So on number 1, the (-0.5) becomes -(0.5) ? @shamil98

OpenStudy (shamil98):

Oh, a typo -(-0.5) = 0.5

OpenStudy (shamil98):

negative x negative = positive so for the first set the product of the multiplied values is positive.

OpenStudy (awkwardpanda):

Thank you both very much (: